Company A is able to sell one of its two milling machines. Both machines perform the same function but differ in age. The newer machine could be sold today for $66,500. Its operating costs are $22,200 a year, but in five years the machine will require a $18,900 overhaul. Thereafter operating costs will be $31,100 until the machine is finally sold in year 10 for $6,650.The older machine could be sold today for $26,100. If it is kept, it will need an immediate $25,500 overhaul. Thereafter operating costs will be $34,000 a year until the machine is finally sold in year 5 for $6,650. Both machines are fully depreciated for tax purposes. The company pays tax at 35%. Cash flows have been forecasted in real terms. The real cost of capital is 15%.What is the equivalent annual costs for selling the new machine?

13974.85

13508.80

70136.55

45283.58
Answer #1

None of the above is the answer

Step 1: Calculate Present Value of Cash Flows from Sale of New Machine

The present value of cash flows that would result from sale of new machine is determined as below:

Present Value = After-Tax Proceeds from Sale of New Machine - Cost of Overhaul of Old Machine - Operating Costs of Old Machine Year 1/(1+Real Cost of Capital)^1 - Operating Costs of Old Machine Year 2/(1+Real Cost of Capital)^2 - Operating Costs of Old Machine Year 3/(1+Real Cost of Capital)^3 - Operating Costs of Old Machine Year 4/(1+Real Cost of Capital)^4 - Operating Costs of Old Machine Year 5/(1+Real Cost of Capital)^5 + After-Tax Proceeds from Sale of Old Machine/(1+Real Cost of Capital)^5

Substituting values in the above formula, we get,

Present Value = 66,500*(1-35%) - 25,500 - 34,000/(1+15%)^1 - 34,000/(1+15%)^2 - 34,000/(1+15%)^3 - 34,000/(1+15%)^4 - 34,000/(1+15%)^5 + 6,650*(1-35%)/(1+15%)^5 = -$94,099.22

_____

Step 2: Calculate Equivalent Annual Cost for Selling the New Machine

The equivalent annual cost for selling the new machine is calculated as below:

Equivalent Annual Cost for Selling New Machine = Present Value of Cash Flows/PVIFA(Rate,Years) = -94,099.22/PVIFA(15%,5) = -94,099.22/3.3522 = $28,071